a proton of mass m and accelerated by a potential difference V gets into a uniform electric field of a parallel plate capacitor parallel to plates of length l at mid point of its separation between the plates. the field strength in it varies with time as E=at, where a is a positive constant. find the angle of deviation of the proton as it comes out of the capacitor.

Grade:12th pass

1 Answers

Nikhil
8 years ago
\theta = \tan^{-1}\left [ al\sqrt{m/32eV^{3}} \right ]    
 
a=positive constant,
l=length of capacitor,
m=mass of proton,
e=1.60218*10^(-19)C,
V=applied potential difference